solve this system with substitution
y=3x+1 and x=3y+1
I don't get it at all and if you could help me out it'd be great:)

Answer by richwmiller(17219)

substitute 3y+1 for x in the first equation
y=3(3y+1)+1
now solve for y
y=9y+3+1
bring all y's to one side all numbers to the other
-4=8y
-4/8=y
-1/2=y
plug that into either original equations
-1/2=3x+1
-3/2=3x
-1/2=x
Yes that is correct
(-1/2,-1/2)